Windsor Company purchased a computer for $8,640 on January 1, 2019. Straight-line depreciation is used, based on a 5-year life and a $1,080 salvage value. On January 1, 2021, the estimates are revised. Windsor now feels the computer will be used until December 31, 2022, when it can be sold for $540.

Compute the 2021 depreciation
Answer #1

Depreciation for 2021 =$2,538

Working

Straight line Method for year 2019 and 2020

A

Cost

$ 8,640

B

Residual Value

$ 1,080

C=A - B

Depreciable base

$ 7,560

D

Life [in years left ]

                                 5

E=C/D

Annual SLM depreciation

$ 1,512

.

Depreciation schedule-Straight line method

Year

Book Value

Depreciation expense

Accumulated Depreciation

Ending Book Value

2019

$ 8,640

$ 1,512

$ 1,512

$ 7,128

2020

$ 7,128

$ 1,512

$ 3,024

$ 5,616

.

Straight line Method -For 2021

A

Ending book value at December 31, 2020

$ 5,616

B

Residual Value

$ 540

C=A - B

Depreciable base

$ 5,076

D

Life [in years left ] (4-2=2)

                                 2

E=C/D

Annual SLM depreciation

$ 2,538